OpenStudy (fluttershyk):

Which graph shows a proportional relationship? (graphs in comments) I think C

OpenStudy (fluttershyk):

@rebeccaxhawaii

rebeccaxhawaii (rebeccaxhawaii):

nope it would not be C

OpenStudy (fluttershyk):

ok then can you explain

rebeccaxhawaii (rebeccaxhawaii):

name the points for me on graph 3

OpenStudy (fluttershyk):

(2,5) (4,8) (6,10)

rebeccaxhawaii (rebeccaxhawaii):

okay so 5/2=2.5 8/4=2 10/6= 1.6666 2.5, 2, and 16666 are not equal pick the one that is

OpenStudy (fluttershyk):

oh its the second one isn't it

rebeccaxhawaii (rebeccaxhawaii):

yes it is good job